Potassium hydrogen phthalate is a solid, monoprotic acid frequently used in the laboratory to standardize strong base solutions. It has the unwieldy formula of KHC8H4O4. This is often written in shorthand notation as KHP. What volume of a 0.163 M calcium hydroxide solution is needed to exactly neutralize 1.37 grams of KHP ?
Answer #1

Ca(OH)2 + 2 KHC8H4O4 = Ca(KC8H4O4)2 + 2 H2O

Moles of KHP = mass / molar mass

= 1.37 / 204.22

= 6.71*10^-3

Moles = Molarity * volume ( in L )

Moles of Ca(OH)2 = 1/2 * moles of KHP ( look balance reaction)

= 1/2 * 6.71 * 10^-3

= 3.355 * 10^-3

0.163 * V = 3.355 * 10^-3

V = 2.06*10^-2 L

Or 20.6 ml

If you have any query please comment

If you satisfied with the solution please rate it thanks